The area of the mural is 33 square meters...what is the radius: The area of the mural is 33 square meters...what is the radius
Answer by checkley77(3848)

area=pir^2
33=3.14*r^2
r^2=33/3.14
r^2=10.51
x=sqrt10.51
x=3.24 meters is the radius.